# Uncertainty in Graph Neural Networks: A Survey

### v1

@@ 1 Introduction
Graph Neural Networks (Gnns) have been widely employed in important real-world applications, including outlier detection (liu2022bond, ), molecular property prediction (wollschlager2023uncertainty, ), and traffic regularization (zhuang2022trafficuncertainty, ). Within these domains, Gnns inevitably present uncertainty towards their predictions, leading to unstable and erroneous prediction results. Uncertainty in Gnns arises from multiple sources, for example, inherent randomness in data, error in Gnn training, and unknown test data from other distributions. To mitigate the adverse impact of predictive uncertainty on Gnn predictions, it is crucial to systematically identify, quantify, and utilize uncertainty. Uncertainty in Gnns plays a vital role in prevalent graph-based tasks, such as selecting nodes to label in graph active learning (cai2017active, ) and indicating the likelihood of a node being out-of-distribution (OOD) (zhao2020uncertainty, ). Furthermore, this practice enhances the predictive performance of Gnns (vashishth2019confidence, ), and bolsters their reliability for decision-making, such as making Gnns robust to attacks (feng2021uag, ) and explains which components of the graphs influence Gnn predictions (ying2019gnnexplainer, ).
@@
Several challenges need to be addressed to boost Gnns with uncertainty. Firstly, different sources of uncertainties in Gnns need to be identified, as uncertainty may stem from multiple stages including data acquisition, Gnn construction, and inference  (gawlikowski2023survey, ). Secondly, different categories of uncertainties should be matched with corresponding tasks. For instance, distributional uncertainty should be applied to graph OOD detection (gui2022good, ). Thirdly, uncertainties in Gnns are difficult to quantify correctly due to the lack of ground truth and unified evaluation metrics. Lastly, quantified uncertainty may be adapted and combined with the other components to better serve the task, e.g., graph active learning (zhang2021alg, ). Multiple studies have endeavored to solve these challenges in specific tasks. In this paper, we provide a holistic view of these tasks through the lens of uncertainty.
@@
To the best of our knowledge, existing surveys or benchmarks either focus on a broader field, i.e., uncertainty in Deep Neural Networks (abdar2021review, ; hullermeier2021aleatoric, ; gawlikowski2023survey, ) and trustworthy Gnn (wu2022survey, ), or on narrower topics, e.g., uncertainty in spatio-temporal Gnns (wu2021trafficquantifying, ; wang2023trafficuncertainty, ). Furthermore, these surveys mainly discuss how to estimate and measure uncertainty, missing concrete and practical guidance on how quantified uncertainties can be applied toward the final downstream task. Our survey aims to bridge these gaps by conducting a systematic review of Gnns from the perspective of uncertainty, with a special focus on incorporating uncertainty into graph learning. Furthermore, our survey contributes to the broader Gnn community by reviewing the trends of the current studies to approach uncertainty from distinct sub-communities, particularly across different downstream tasks. Our goal is to present a comprehensive framework that clarifies ambiguous terminologies, connects several communities, as well as offers valuable insights and directions to advance the overall development of Gnn uncertainty and related fields.
@@
In this paper, we organize the literature of uncertainty in Gnns into a three-stage framework, illustrated in Figure 1. Consequently, we introduce real-world applications utilizing Gnn uncertainty in Section 5. Finally, in Section 6, we discuss significant and promising avenues for future research in this field.
@@
Figure 1: Overall Framework: (1) identifying sources of uncertainty (Section 2), (2) quantifying uncertainty (Section 3) and (3) utilizing uncertainty for downstream tasks (Section 4).
@@ 2 Identifying Sources of Uncertainty in Gnns
In machine learning, uncertainty generally equals predictive uncertainty (PU), defined as the uncertainty propagated into the prediction results. For accurate estimation and application of uncertainty, it is necessary to first identify different sources of uncertainty. Predictive uncertainty is commonly divided into aleatoric and epistemic ones based on their inherent different sources (hullermeier2021aleatoric, ).
@@
Aleatoric uncertainty (AU). Aleatoric uncertainty—known as data or statistical uncertainty—refers to variability in the outcome of an experiment which is due to inherent random effects. This often arises in the pre-processing stage, where the data collection may lack sufficient information, or contain noise and error. For example, a graph outlier detection dataset may fail to label all abnormal nodes due to the limited label cost. Notably, different from munikoti2023general , we follow the mainstream theory that aleatoric uncertainty can only be estimated but remains non-reducible by improving the Gnn (hullermeier2021aleatoric, ).
@@
Epistemic uncertainty (EU). In contrast, epistemic uncertainty—known as systemic uncertainty—refers to the uncertainty caused by the lack of knowledge of the Gnn, thus reducible. It arises in the procedure of Gnn modeling, further classified as model uncertainty (MU). On one hand, model uncertainty comes from the choice of model structure. For instance, the shallow network structure may lead to the under-confidence of Gnns (wang2022gcl, ). On the other hand, the training process of Gnn also introduces model uncertainty in the selection of parameters, for example, learning rate, epoch number, batch size, and techniques (e.g., weight initialization). Another type of epistemic uncertainty is distributional uncertainty (DU), introduced by malinin2018predictive . While applying the trained Gnn to the test data for inference, the distribution shift contributes to the uncertainty of prediction, such that the training and test sets do not come from the same distribution. This is common for inductive inference, where Gnn is trained on one graph and applied to another graph for prediction tasks.
@@
Besides these frequently classified types, Gnn uncertainty can also be classified from subjective viewpoints, e.g., vacuity (conflicting evidence) and dissonance (lack of evidence) (zhao2020uncertainty, ). In general, the goal for Gnn modeling should be to reduce EU, including delicate design of model structures and parameters, along with enhanced robustness and generalization to diverse distributions. On the contrary, since AU can only be quantified but not reduced, separating it from the epistemic uncertainty would provide a more reliable illustration of Gnn performance.
@@ 3 Quantifying Uncertainty in Gnns
In this section, we classify existing Gnn uncertainty quantification methods into three classes according to the number of models as well as whether the model parameters are random or deterministic, followed by evaluation metrics for these methods. In Table 1, we outline a summary of various quantification methods.
@@
Properties Single Deterministic Model Single Model with Random Parameters Ensemble Model
@@
Cost of Training Low High High
@@ 3.1 Single Deterministic Model
Direct Estimation. The most commonly used uncertainty estimation method is through the softmax probability of prediction outcomes, including maximum softmax probability (wang2021confident, ) and entropy (cai2017active, ). Direct estimation methods integrate uncertainty estimation with prediction and do not require modifications of the original model.
@@
It has been shown that Gnn tends to be under-confident (wang2021confident, ). Confidence is an indication of how likely (probability) the predictions of a machine learning algorithm are correct, often computed as the maximum softmax class probability. Under-confidence, in other words, refers to that confidence, as the indication of accuracy, falls below the true prediction accuracy. To mitigate the issue of under-confidence, a well-calibrated Gnn should produce predicted probabilities that accurately reflect the true likelihood of the predicted outcomes. This is usually evaluated with a smaller Expected Calibration Error (ECE) (guo2017calibration, ). To provide faithful uncertainty estimation, wang2021confident add a regularizer in the Gnn optimization objective, which penalizes the probabilities of the incorrect nodes and encourages the correct nodes to be more confident. liu2022calibration employ classical post-processing methods to calibrate Gnn confidence, including histogram binning, isotonic regression, and Bayesian binning into quantiles. Similarly, hsu2022makes propose a post hoc calibration method tailored for Gnns.
However, the inherent problem with estimated uncertainty is that softmax output only reflects the total predictive uncertainty instead of the model uncertainty, leading to false confidence under the distribution shift. Other heuristic measures, such as the distance of the sample to each class in the representation space (xu2023open, ), are also used to quantify uncertainty.
Bayesian-based Estimation. Another family of Gnn uncertainty estimation methods is the Bayesian-based deterministic model. In contrast to direct estimation, the model fulfills the dual objectives of prediction and uncertainty estimation using Bayesian statistical principles. These methods typically utilize the Dirichlet distribution, which serves as the conjugate prior distribution of the categorical distribution in Bayesian inference.
@@
Dirichlet distribution is advantageous in directly reflecting higher distributional uncertainty with flatter (less variant) logits over all categories. Therefore, it is utilized in Gnn uncertainty estimation approaches such as Evidential neural networks (zhao2020uncertainty, ; stadler2021graph, ; bazhenov2023revisiting, ). zhao2020uncertainty first apply Dirichlet-based methods to Gnns, which incorporates information about the local neighborhood through the shortest path distances and the node labels. stadler2021graph better model the network effects through diffusing Dirichlet parameters in graph propagation.
As these methods are also trained to quantify uncertainty, they are more suitable for specific tasks, e.g., OOD detection, but may not produce the best performance in other downstream tasks, e.g., node classification (bazhenov2023revisiting, ).
@@
Frequentist-based estimation is a recently adverted family of methods to achieve trustworthy uncertainty estimation. Frequentist methods are typically model-agnostic and estimate uncertainty through post-processing. This implies that they do not alter the model or directly offer uncertainty estimation alongside prediction.
@@
In contrast to Bayesian-based approaches, they exhibit computational efficiency with fewer additional parameters, and their estimation is more robust compared to the direct methods. kang2022jurygcn estimate uncertainty through the length of confidence bands, constructed with jackknife (leave-one-out) sampling (efron1982jackknife, ) on training data and influence function to estimate the change in Gnn parameters. Though this idea is similar to MC dropout in Section 3.2, it leaves out samples and modifies model parameters in a Frequentist manner instead of Bayesian approaches, which assume a prior distribution of model components. Employing the conformal prediction framework, huang2023uncertainty and zargarbashi2023conformal provide guaranteed uncertainty estimation with the size of prediction sets. Though both methods optimize for smaller set sizes (i.e., better efficiency) while covering the true labels in sets with guaranteed high probability, zargarbashi2023conformal propagate node labels in the calibration set, and huang2023uncertainty directly use the inefficiency of the calibration set as the optimization objective.
While Frequentist-based methods are more robust and may adapt to distribution shifts, the existing ones do not explicitly distinguish between AU and EU.
3.2 Single Model with Random Parameters
@@
Models with random parameters, mostly Bayesian Gnns, are another widely used family of Gnn uncertainty quantification methods. A vast majority of these methods rely on Monte Carlo (MC) dropout to sample weights for variational inference. zhang2019bayesian first develop a Bayesian neural network framework for graphs depending on the graph topology, where the observed graph is viewed as a realization from a parametric family of random graphs. hasanzadeh2020bayesian point out that the graph dropout techniques (rong2019dropedge, ) can be utilized at the test time for uncertainty estimation, and advance zhang2019bayesian by considering node features and adaptive connection sampling. cha2023temperature discover that under the conformal prediction framework, Bayesian Gnn with a tempered posterior can improve the efficiency of predictions. Markov Chain Monte Carlo (MCMC) sampling methods are another family of Bayesian methods that represent uncertainty without a parametric model. However, these models are rarely used in graphs due to their expensive computational costs (wu2021trafficquantifying, ).

### v2

@@ 1 Introduction
Graph Neural Networks (Gnns) have been widely employed in important real-world applications, including outlier detection (Liu et al., 2022b), molecular property prediction (Wollschläger et al., 2023), and traffic regularization (Zhuang et al., 2022). Within these domains, Gnns inevitably present uncertainty towards their predictions, leading to unstable and erroneous prediction results. Uncertainty in Gnns arises from multiple sources, for example, inherent randomness in data, error in Gnn training, and unknown test data from other distributions. To mitigate the adverse impact of predictive uncertainty on Gnn predictions, it is crucial to systematically identify, quantify, and utilize uncertainty. Uncertainty in Gnns plays a vital role in prevalent graph-based tasks, such as selecting nodes to label in graph active learning (Cai et al., 2017) and indicating the likelihood of a node being out-of-distribution (OOD) (Zhao et al., 2020). Furthermore, this practice enhances the predictive performance of Gnns (Vashishth et al., 2019), and bolsters their reliability for decision-making, such as making Gnns robust to attacks (Feng et al., 2021) and explains which components of the graphs influence Gnn predictions (Ying et al., 2019).
@@
Several challenges need to be addressed to boost Gnns with uncertainty. Firstly, different sources of uncertainties in Gnns need to be identified, as uncertainty may stem from multiple stages including data acquisition, Gnn construction, and inference  (Gawlikowski et al., 2023). Secondly, different categories of uncertainties should be matched with corresponding tasks. For instance, distributional uncertainty should be applied to graph OOD detection (Gui et al., 2022). Thirdly, uncertainties in Gnns are difficult to quantify correctly due to the lack of ground truth and unified evaluation metrics. Lastly, quantified uncertainty may be adapted and combined with the other components to better serve the task, e.g., graph active learning (Zhang et al., 2021a). Multiple studies have endeavored to solve these challenges in specific tasks. In this paper, we provide a holistic view of these tasks through the lens of uncertainty.
@@
To the best of our knowledge, existing surveys or benchmarks either focus on a broader field, i.e., uncertainty in Deep Neural Networks (Abdar et al., 2021; Hüllermeier & Waegeman, 2021; Gawlikowski et al., 2023) and trustworthy Gnn (Wu et al., 2022), or on narrower topics, e.g., uncertainty in spatio-temporal Gnns (Wu et al., 2021; Wang et al., 2023). Besides, these surveys mainly discuss how to estimate and measure uncertainty, missing concrete and practical guidance on how quantified uncertainties can be applied toward the final downstream task. Our survey aims to bridge these gaps by conducting a systematic review of Gnns from the perspective of uncertainty, with a special focus on incorporating uncertainty into graph learning. Furthermore, our survey contributes to the broader Gnn community by reviewing the trends of the current studies to approach uncertainty from distinct sub-communities, particularly across different downstream tasks. Our goal is to present a comprehensive framework that clarifies ambiguous terminologies, connects several communities, as well as offers valuable insights and directions to advance the overall development of Gnn uncertainty and related fields.
@@
In this paper, we organize the literature of uncertainty in Gnns into a three-stage framework, illustrated in Figure 1. Consequently, we introduce real-world applications utilizing Gnn uncertainty in Section 5. Finally, in Section 6, we discuss significant and promising avenues for future research in this field. For further related background, we encourage readers to investigate Wu et al. (2020b) for a detailed introduction to Gnns, and Abdar et al. (2021); Hüllermeier & Waegeman (2021); Gawlikowski et al. (2023) for uncertainty in deep learning.
@@
Figure 1: Overall Framework: (1) identifying sources of uncertainty (Section 2), (2) quantifying uncertainty (Section 3) and (3) utilizing uncertainty for downstream tasks (Section 4). In the first subfigure, ‘PU’ refers to predictive/total uncertainty. ’AU’ represents aleatoric/data/statistical uncertainty, ’EU’ is in short for epistemic uncertainty. ’MU’ and ’DU’ represent model and distributional uncertainty, respectively.
@@ 2 Identifying Sources of Uncertainty in Gnns
In machine learning, (total) uncertainty generally equals predictive uncertainty (PU), defined as the uncertainty propagated into the prediction results. For accurate estimation and application of uncertainty, it is necessary to first identify different sources of uncertainty. Predictive uncertainty is commonly divided into aleatoric and epistemic ones based on their inherent different sources (Hüllermeier & Waegeman, 2021).
@@
Aleatoric uncertainty (AU). Aleatoric uncertainty—known as data or statistical uncertainty—refers to variability in the outcome of an experiment which is due to inherent random effects. This often arises during the pre-processing stage, where data collection may lack sufficient information or contain noise and errors. As an example, collecting more data from the same graph distribution does not reduce aleatoric uncertainty; however, incorporating additional node features with non-overlapping information can reduce the inherent randomness. Unlike Munikoti et al. (2023), we align with the prevailing theory that aleatoric uncertainty can be estimated but is non-reducible through improvements in Gnn (Hüllermeier & Waegeman, 2021), using it as the primary criterion to categorize existing literature.
@@
Epistemic uncertainty (EU). In contrast, epistemic uncertainty—known as systemic uncertainty—refers to the uncertainty caused by the lack of knowledge of the Gnn, thus reducible. It arises in the procedure of Gnn modeling, further classified as model uncertainty (MU). On one hand, model uncertainty comes from the choice of model structure. For instance, the shallow network structure may lead to the under-confidence of Gnns (Wang et al., 2022). On the other hand, the training process of Gnn also introduces model uncertainty in the selection of parameters, for example, learning rate, epoch number, batch size, and techniques (e.g., weight initialization). Introduced by Malinin & Gales (2018), another type of epistemic uncertainty is distributional uncertainty (DU), which arises when Gnn has false assumption about the problem setting and data generation process. This is common for inductive inference, where Gnn is trained on one graph and applied to another graph for prediction tasks, such that the training and test sets do not come from the same distribution. While applying the trained Gnn to the test data for inference, the distribution shift contributes to the uncertainty of prediction.
Note that since EU = MU + DU, and PU = AU + EU, then PU = AU + MU + DU. This relationship is illustrated in the middle subfigure of Figure 2. In most settings, prediction results are typically considered to have no distributional uncertainty, DU=0, thus EU=MU+0=MU. To avoid confusion of terms, we will use epistemic uncertainty (EU) at most time in Sections 4 and 5. If distributional uncertainty (DU) is explicitly stated, such as in tasks like OOD, we will use MU and DU.
Besides these frequently classified types, Gnn uncertainty can also be classified from subjective viewpoints, e.g., vacuity (conflicting evidence) and dissonance (lack of evidence) (Zhao et al., 2020).
@@
In general, the goal for Gnn modeling should be to reduce EU, including delicate design of model structures and parameters, along with enhanced robustness and generalization to diverse distributions. On the contrary, since AU can only be quantified but not reduced, separating it from the epistemic uncertainty would provide a more reliable illustration of Gnn performance.
@@ 3 Quantifying Uncertainty in Gnns
In this section, we classify existing Gnn uncertainty quantification methods into three classes according to the number of models and whether model parameters are random variables. Following this classification, we present evaluation metrics for different quantification methods. In Table 1, we summarize and compare attributes of these methods.
@@
Properties Single Deterministic Model Single Model with Parameters Being Random Variables Ensemble Model
@@
Cost of Training Low Higher High
@@ 3.1 Single Deterministic Model
Direct estimation. The most commonly used uncertainty estimation method is through the softmax probability of prediction outcomes, including maximum softmax probability (Wang et al., 2021) and entropy (Cai et al., 2017). Other heuristic measures, such as the distance of the sample to each class in the representation space (Xu et al., 2023), are also used to quantify uncertainty. Direct estimation methods integrate uncertainty estimation with prediction and do not require modifications of the original model.
@@
However, it has been shown that Gnn tends to be under-confident (Wang et al., 2021), which leads to inaccurate estimation of confidence and entropy. Confidence is an indication of how likely (probability) the predictions of a machine learning algorithm are correct, often computed as the maximum softmax class probability. Under-confidence, in other words, refers to that confidence, as the indication of accuracy, falls below the true prediction accuracy. To mitigate the issue of under-confidence, a well-calibrated Gnn should produce predicted probabilities that accurately reflect the true likelihood of the predicted outcomes. This is usually evaluated with a smaller Expected Calibration Error (ECE) (Guo et al., 2017). To provide faithful uncertainty estimation, Wang et al. (2021) add a regularizer in the Gnn optimization objective, which penalizes the probabilities of the incorrect nodes and encourages the correct nodes to be more confident. Liu et al. (2022d) employ classical post-processing methods to calibrate Gnn confidence, including histogram binning, isotonic regression, and Bayesian binning into quantiles. Similarly, Hsu et al. (2022b) propose a post hoc calibration method tailored for Gnns. Compared with uncalibrated direct estimation, the current calibration methods can effectively reduces ECE, but require additional calibration samples.
In addition, the inherent problem with estimated uncertainty is that the softmax output only reflects the total predictive uncertainty instead of the model uncertainty, leading to false estimation under the distribution shift. Therefore, direct estimation is a simple and easy-to-apply choice only when computation and data efficiency is preferable over estimation accuracy.
@@
Bayesian-based estimation is a family of Bayesian-based deterministic models. Compared to direct estimation, the model fulfills the dual objectives of prediction and uncertainty estimation using Bayesian statistical principles. These methods typically utilize the Dirichlet distribution, which serves as the conjugate prior distribution of the categorical distribution in Bayesian inference. Dirichlet distribution is advantageous in directly reflecting higher distributional uncertainty with flatter (less variant) logits over all categories. Therefore, it is utilized in Gnn uncertainty estimation approaches such as Evidential neural networks (Zhao et al., 2020; Stadler et al., 2021; Bazhenov et al., 2023). Zhao et al. (2020) first apply Dirichlet-based methods to Gnns, which incorporates information about the local neighborhood through the shortest path distances and the node labels. Stadler et al. (2021) better model the network effects through diffusing Dirichlet parameters in graph propagation. As the training loss are modified for both prediction accuracy and uncertainty quantification, especially for DU, these methods are more suitable for specific tasks, e.g., OOD detection, but may not produce the best performance in other tasks, e.g., node classification (Bazhenov et al., 2023). For the same reason, another disadvantage of existing methods is that they can not be directly applied on already trained Gnns post-hoc.
@@
Frequentist-based estimation. This is a recently adverted family of methods to achieve trustworthy uncertainty estimation. Frequentist methods are typically model-agnostic and estimate uncertainty through post-processing. This implies that they do not alter the model or directly offer uncertainty estimation alongside prediction. In contrast to Bayesian-based approaches, they exhibit computational efficiency with fewer additional parameters, and their estimation is distribution-free and more robust compared to the direct methods.
@@
Kang et al. (2022) estimate uncertainty through the length of confidence bands, constructed with jackknife (leave-one-out) sampling (Efron, 1982) on training data and influence function to estimate the change in Gnn parameters. Though this idea is similar to MC dropout in Section 3.2, it leaves out samples and modifies model parameters in a Frequentist manner, not assuming a prior distribution of model components. Early works (Bååw, 2022; Clarkson, 2023) employing conformal prediction framework to Gnns provide guaranteed uncertainty estimation with the size of prediction sets under different settings, but involve less specific designs for Gnns. Recent works (Huang et al., 2023a; Zargarbashi et al., 2023) focus on optimizing for smaller set sizes (i.e., better efficiency) while covering the true labels in sets with guaranteed high probability. To achieve this goal, Zargarbashi et al. (2023) propagate node labels in the calibration set, and Huang et al. (2023a) directly use the inefficiency of the calibration set as the optimization objective.
However, existing Frequentist-based methods quantify total predictive uncertainty and do not explicitly distinguish between AU and EU, as opposed to the belief that they only quantify AU (Qian et al., 2023). It is also noteworthy that all these methods are built on the data exchangeability assumption. To satisfy this assumption, current methods split the data in fixed modes, mostly via random sampling. Lunde et al. (2023) provide some other general sampling methods in graphs where exchangeability assumption is not violated. However, in these cases, some specific graph properties are required. The validity of methods using non-random data splits, e.g., inductive settings, rely on nonexchangeable conformal prediction techniques, especially weighted conformal prediction (Barber et al., 2023). This area of research still has ample space for further exploration, particularly regarding how to optimally choose the weights. In practice, exchangeability assumption is believed to be satisfied if the coverage rate meets the pre-determined level in the experiments. Though this is not a sufficient condition to support the assumption, this practice is reasonable as guaranteed coverage rate is part of the method’s goal and evaluation metric. Furthermore, in current graph conformal prediction methods, the guarantee of uncertainty quantification is marginal coverage on the total test data or conditional on some specific groups. Conditioning on each sample requires uniform unconformity scores (Einbinder et al., 2022), which are nearly impossible. This hinders the application of conformal methods on most downstream tasks in Section 4.
3.2 Single Model with Parameters Being Random Variables
@@
Models with parameters being random variables, mostly Bayesian Gnns, are another widely used family of Gnn uncertainty quantification methods. A vast majority of these methods rely on Monte Carlo (MC) dropout to sample weights for variational inference. Zhang et al. (2019b) first develop a Bayesian neural network framework for graphs depending on the graph topology, where the observed graph is viewed as a realization from a parametric family of random graphs. Hasanzadeh et al. (2020) point out that the graph dropout techniques (Rong et al., 2019) can be utilized at the test time for uncertainty estimation, and advance Zhang et al. (2019b) by considering node features and adaptive connection sampling. Cha et al. (2023) discover that under the conformal prediction framework, Bayesian Gnn with a tempered posterior can improve the efficiency of predictions. Markov Chain Monte Carlo (MCMC) sampling methods are another family of Bayesian methods that represent uncertainty without a parametric model. However, these models are rarely used in graphs due to their expensive computational costs (Wu et al., 2021).